OpenStudy (anonymous):

Graph the following function using transformations. Find the vertical and horizontal asymptotes. State the domain and range. f(x) = 5+ 1/x+4. Please show work. @Calculus1

OpenStudy (aravindg):

i advice u to use wolfralpha

OpenStudy (anonymous):

is this like this? f(x) = 5+ 1/(x+4)

OpenStudy (anonymous):

MArk o...it doesnt show parenthesis around...

OpenStudy (agreene):

so, is: \[f(x)=5+\frac{1}{x}+4\] correct? or is mark's interpretation correct: \[f(x)=5+\frac{1}{x+4}\] ?

OpenStudy (anonymous):

the second one...agreene...

OpenStudy (agreene):

Okies. For the vertical asymptotes, you need to find areas where the function wont be defined. in this case where null division happens. x+4=0 x=-4 So, we have a vertical asymptote at x=-4 For the horizontal, we then look to the highest powers of x, and ignore everything else giving us: \[\frac{1}{x}\] we then take away the x's giving us our horizontal asymptote: \[\frac{0}{1}=0\] So, our asymptotic regions are: x=-4 (Vertical) and x=0 (Horizontal) As for domain and range, there are no restricting agents in the function, causing \[-\infty\le x\le \infty\] \[-\infty \le y \le \infty\]

OpenStudy (anonymous):

Thank you agreene!

OpenStudy (anonymous):

agreene- so what would the graph look like then??